OpenStudy (anonymous):

HELP PLEASE!!:c Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12? x greater than or equal to -5 x less than or greater to -5 -5 less than or greater to x A number line with a closed circle on negative 5 and shading to the right

OpenStudy (texaschic101):

2(x - 1) >= -12 -- distribute through the parenthesis 2x - 2 >= -12 -- add 2 to both sides 2x >= -12 + 2 2x >= -10 x >= -5 This is stating that x is greater or equal to -5. Or, -5 is less then or equal to x. It will have a closed circle on -5, shading to the right. So what do you think ?

OpenStudy (anonymous):

Thank you, Thank you sooo much! c:

OpenStudy (texaschic101):

anytime :)